Please adhere to the rounding instructions(in red) for calculations and final answer. If the answer is not exact, it will be incorrect. Thanks!

Question 1 0.5 pts What is the future value in 7.8 years of $5546 invested in an account with a stated annual interest rate of 8%, compounded annually? (Round answer to 2 decimal places. Do not round intermediate calculations). Topic: Discounted Cash Flow Valuation
Answer #1

FV= PV(1+ R/(n*100))^n

R= annual rate of interest=0.08

n= number of compounding periods i.e 1 in our case because compounded annually

FV= Future value

PV= Present Value i.e 5546

therefore:

FV= 5546*(1+0.08)^7.8 = 10,108.46